# Principled priors for Bayesian inference of circular models

### v2

@@ 4.3 PC Prior for Wrapped Cauchy Distribution
$\pi\left(\rho\right)$ $=\lambda\exp\left\{-\lambda\sqrt{-\log\left(1-\rho^{2}\right)}\right\}\frac{\rho}{\left(1-\rho^{2}\right)\sqrt{-\log\left(1-\rho^{2}\right)}},$ (28)
@@
The proof is provided in Section A.3, and the density and log-log density plots of this PC prior are shown in Figure 8. The density plot indicates that higher values of $\lambda$ result in greater prior density at $\rho=0$, and the prior density increases sharply as $\rho$ approaches 1. In the log-log scale, the prior behavior reveals a consistent trend across different values of $\lambda$, suggesting similar characteristics despite variations in $\lambda$.
Figure 8: PC prior density ($\pi\left(\rho\right)$, left) and log-log density ($\log\pi\left(\log\rho\right)$, right).
@@ 5 Simulation Study and Comparisons
The logic behind our proposed prior selection framework lies in checking whether a prior favors simpler models to prevent overfitting. To further illustrate the way of using our framework, and to demonstrate the advantages of employing PC priors for circular models, this section compares the proposed priors in Section 4.1 with existing priors from the literature. The comparison focuses on their behavior in both the parameter scale and the distance scale (Section 5.1). Additionally, a comprehensive set of simulation studies is presented in Section 5.2.
5.1 Properties of the PC and Other Common Priors
Recall the procedure provided in Section 3.3, the PC prior is exponential in distance scale and therefore inherently favors simpler models and helps prevent overfitting. Figure 9 shows the PC prior density in distance scale. Since the PC prior always performs as an exponential distribution in distance scale (Equation 9), it always assigns maximum density at $d=0$. It can be seen that the larger the value of scaling parameter $\lambda$ is, the more heavily the prior density will be penalized when the distance increases.
Figure 9: The PC prior density in distance scale.
To check the distance scale behavior of other priors, the transformation in Equation 8 is required.
@@ 5.1.1 Von Mises Distribution
Figure 10 illustrates the behaviors of the PC prior, the popular $\operatorname{Gamma}(1,b)$ prior, $h_{2}$ prior (Equation 3), and $h_{3}$ prior (Equation 3) in both the parameter scale and the distance scale. The distance measure is defined between the von Mises (vM) density and the circular uniform density. For fairness, the scaling parameter $\lambda$ of the PC prior and the shape parameter $b$ of the $\operatorname{Gamma}(1,b)$ prior are calibrated so that both priors share the same median.
From the plot of the density on the distance scale, it is evident that the $h_{3}$ prior assigns zero density at $d=0$, indicating that it never suggests the data being uniform. This behavior leads to overfitting issues. The $\operatorname{Gamma}(1,b)$ prior has reasonable density at the base model, whereas it exhibits a non-monotonic density with a global maximum around $d\approx 1$ rather than at $d=0$, which indicates poor robustness in preventing overfitting. The $h_{2}$ prior does not penalize the density exponentially with respect to the distance by nature, however, it has a reasonable behavior with respect to distance. Its density peaks at $d=0$ and decreases monotonically as $d\to\infty$.
Figure 10: Priors for $\kappa$ in parameter (left) and distance scale with the base model at $\kappa_{0}=0$ (right). The solid line is the $\operatorname{PC}\left(\lambda=0.92\right)$ prior, the dashed line is the $\operatorname{Gamma}\left(1,0.34\right)$, the dotted line is the $h_{2}$ prior and the dot-dash line is the $h_{3}$ prior.
@@ 5.1.2 Cardioid Distribution
Figure 13 and Figure 13 show the boundary behaviors for the priors for parameter $\ell$ with the base models being at $\ell_{0}=0$ and $\ell_{0}\rightarrow 0.5$ respectively. The plots conclude that only PC priors perform well at the boundaries, assigning non-zero density at $d=0$, whilst the density of uniform and $2\times\operatorname{Beta}$ priors are not even smooth. This feature illustrates the possible inherent caveat when priors are specified to be smooth and well-behaved on the parameter scale.
@@ 5.1.3 Wrapped Cauchy Distribution
The Beta prior is one of the most commonly used choices for the concentration parameter $\rho$ of the wrapped Cauchy distribution. The prior density on both the parameter scale and the distance scale is illustrated in Figure 14. The behavior of the Beta prior is particularly interesting, as it exhibits different trends near the two boundaries depending on the choice of the $a$ and $b$ hyperparameters. The right-hand plot shows that the Beta prior can avoid overfitting issues only when $a<0$, ensuring the density at $d=0$ is non-zero.
Figure 14: Beta prior and PC ($\lambda=1$) prior density in parameter ($\pi\left(\rho\right)$, left) and distance scale ($\pi\left(d\right)$, right).
Therefore, a Beta prior with $a<1$ is considered acceptable. When users have sufficient information to set an informative prior, a Beta prior with $a<0$ is a reasonable choice, as it maintains the ability to prevent “overconfidence”. However, selection of the hyperparameter $b$ significantly influences the behavior of the Beta prior. Thus, we should always be careful when employing this beta prior in practice.
@@ 5.2 Simulation Study
In former subsections, we illustrate the way of employing our proposed framework, and the behavior of the discussed priors are illustrated and examined. In this section, we further show the performance of the priors and the robustness of the proposed circular PC priors through comprehensive simulation studies.
The simulation study is conducted through comparing the posteriors of the concentration parameter $\xi$ ($\kappa$ for von Mises distribution, $\ell$ for cardioid distribution and $\rho$ for wrapped Cauchy distribution) under different priors. For consistency, $\pi\left(\mu\right)$ is chosen to be the circular uniform density, thus, the joint posterior is as follows:
$\pi\left(\mu,\xi\mid x\right)$ $\propto\pi\left(x\mid\mu,\xi\right)\pi\left(\mu\right)\pi\left(\kappa\right).$ (31)
In the study, posteriors are derived from samples according to the HMC framework available in the STAN interface (Carpenter et al., 2017), and the code can be found in https://github.com/XiangYEstats/circular_PCpriors_stan. For consistency, the data are sampled from $\pi\left(x\mid\mu,\xi\right)$ (vM, cardioid and WC distribution respectively) with $N=100,300,1000$ sample sizes and location parameter $\mu=\pi$. In addition, the result is obtained by taking the mean of 100 replicates of the experiment with all random seeds set to be 520 equally.
@@ 5.2.1 Von Mises Distribution
For $\pi\left(\kappa\right)$ we consider the $\operatorname{Gamma}\left(1,b\right)$ priors, the $h_{2}$ and the $h_{3}$ prior (Equation 3), and the PC priors with uniform base model (PCU, Equation 12) and point mass base model (PCP, Equation 14). The setting for the simulation study for $\kappa$ is given below:
  • The data are sampled with $\kappa=0.02,0.33,1,1.67,3,7,15,59$;
  • The hyperparameter for $\operatorname{Gamma}\left(1,b\right)$ prior has values of $b=0.01,0.05,0.1,1,5$;
  • The scaling parameters for both PCU and PCP priors are set with $U=\frac{\pi}{2}$, and $\alpha=0.01,0.1,0.3,0.5,0.7,0.9,0.99$.
From the results in Figure 15, we can see that both PCU and PCP priors perform robustly well with different values of $\alpha$ (resulting in different values of $\lambda$), whilst the posterior for $\operatorname{Gamma}\left(1,b\right)$ priors act differently especially when the sample size is small and the concentration is high. This fact indicates that the $\operatorname{Gamma}\left(1,b\right)$ prior highly relies on the value of hyperparameter, wheares the PC priors perform well with different values of $U$ and $\alpha$.
Notably, the $h_{2}$ prior also performs well in fitting the model. However, based on the discussion in Section 5.1.1, it is not recommended to employ $h_{2}$ prior when the data are highly concentrated, as it will overfit the model.
Figure 15: Simulation Study for the priors for concentration parameter $\kappa$ of von Mises distribution.
@@ 5.2.2 Cardioid Distribution
For $\pi\left(\ell\right)$ we consider the $\operatorname{Uniform}\left(0,0.5\right)$ prior, the $2\times\operatorname{Beta}\left(a,b\right)$ prior and the PC priors with the uniform base model (PCU, Equation 19) and the cardioid curve base model (PCC, Equation 22). The setting for the simulation study for $\ell$ is given below:
  • The data are sampled with $\ell=0,0.01,0.1,0.2,0.3,0.4,0.49$;
  • The hyperparameter for $2\times\operatorname{Beta}\left(a,b\right)$ prior has values of $a,b\in\left\{0.5,1,2,5\right\}$;
  • The scaling parameters for both PCC and PCU priors are set with $U=0.5$, but with $\alpha=0.01,0.1,0.3,0.5,0.7,0.9,0.99$ for PCC prior and $\alpha=0.01,0.1,0.2,0.3,0.4,0.5$ for PCU prior, since $\lambda$ takes very small values when $\alpha>0.5$ for PCU prior.
Figure 16: Simulation Study for the priors for concentration parameter $\ell$ of cardioid distribution.
From the results given in Figure 16, we could observe that both PCU and PCC priors have much more stable and accurate performance than the $2\times\operatorname{Beta}\left(a,b\right)$ prior. Moreover, the PCU prior performs better than the PCC prior around the $\ell=0$ boundary, whilst the PCC prior naturally perform better around the $\ell\rightarrow 0.5$ boundary. This fact demonstrates the importance of base model selection when employing PC prior. In addition, the $\operatorname{Uniform}\left(0,0.5\right)$ prior performs well in the simulation study except for the situation of being at the $\ell=0$ boundary. However, as discussed in Section 5.1.2, the $\operatorname{Uniform}\left(0,0.5\right)$ prior is also not recommended as it will also overfit the data.
@@ 5.2.3 Wrapped Cauchy Distribution
For $\pi\left(\rho\right)$ we consider the $\operatorname{Beta}\left(a,b\right)$ and the PC priors (Equation 28). The setting for this study is given below:
  • The data are sampled with $\rho=0,0.01,0.1,0.3,0.5,0.7,0.9,0.99$;
  • The hyperparameter for $\operatorname{Beta}\left(a,b\right)$ prior has values of $a,b\in\left\{0.5,1,2,5\right\}$;
  • The scaling parameters for PC prior are set with $U=0.6$, and $\alpha=0.01,0.1,0.3,0.5,0.7,0.9,0.99$.
Figure 17: Simulation Study for the priors for concentration parameter $\rho$ of wrapped Cauchy distribution.
Figure 17 shows that the PC priors demonstrate more stable and accurate performance than the $\operatorname{Beta}\left(a,b\right)$ prior. However, if there exists sufficient information supporting a $\operatorname{Beta}\left(a,b\right)$ prior, as discussed in Section 5.1.3, we suggest that the $\operatorname{Beta}\left(a,b\right)$ prior with $a<1$ are proper choices since they would not lead to overfitting issue.
@@ 6 Application
In this section, we present a real-data case study to show the practical performance of the discussed priors, and the code can be found in https://github.com/XiangYEstats/circular_PCpriors_stan. The data used in this study are the wind data (Figure 20) stored in “circular” package in R, recorded by a meteorological station in a place named “Col de la Roa” in the Italian Alps. The dataset contains 310 measures of daily wind direction from January 29th, 2001 to March 31st, 2001 covering the data span from 3 a.m. to 4 a.m. of the day. As a circular analogue of normal distribution, von Mises distribution is employed for our model, and the Bayesian model is the same as expressed in Equation 31. For comparison, the Stan MCMC model for Equation 31 with $\pi\left(\kappa\right)$ being PC prior, $\operatorname{Gamma}\left(1,b\right)$ prior (exponential prior), $h_{2}$ prior and $h_{3}$ prior for $\kappa$ are constructed.
We choose $b=0.1,1,2,5,10$ for the $\operatorname{Gamma}\left(1,b\right)$ prior. From Figure 20, we can observe that this dataset is not highly concentrated. Therefore, for the PC prior, the circular uniform base model would be a more natural choice compared with the point mass base model. However, the purpose of comparison and illustration, models with PC prior with both base models are constructed. We choose the scaling parameter $\lambda$ following the procedure discussed in Section 3.4 with $U=\pi,\frac{\pi}{2},\frac{\pi}{4},\frac{\pi}{8},\frac{\pi}{16}$, indicating different radians around $x=0$. The $\alpha$ value is computed by finding ratio between the number of data lying in the chosen radian $U$ and 310 (total amount of data). An example code is given below as an illustration:
Figure 18: Example code for choosing $\alpha$ values.
After fitting the model, the posterior mean and the 95% credible interval for $\kappa$ are given in Figure 20. From the plot, it is obvious that both the PC prior with circular uniform base model (PCU) and PC prior with point mass base model (PCP) fit the data well with the proposed scaling parameter choosing strategy. For this dataset, as a more natural choice, the PCU prior is indeed robust than PCP prior, since the value of scaling parameter $\lambda$ does not vary much with the same given $U$ and $\alpha$ values. However, the PCP prior still shows its robustness, compared with the $\operatorname{Gamma}\left(1,b\right)$ prior, as the most extreme values for $\lambda$ in our example still lead to smaller errors in posterior, and, moreover, there is insufficient guidance for choosing the values for hyperparameters of $\operatorname{Gamma}\left(1,b\right)$ prior. In addition, the $h_{2}$ and $h_{3}$ priors both perform well in this example. In Section 5.1.1, we discussed that $h_{2}$ prior also performs well in preventing overfitting when the base model is the circular uniform distribution. Thus, to prevent overfitting, the $h_{2}$ prior also appears to be a proper choice of prior for $\kappa$ when the data appear to have low concentration.
@@
Figure 20: Posterior means and 95% credible intervals of $\kappa$.

### v3

@@ 3.2.3 Wrapped Cauchy Distribution
$p\left(\rho\right)$ $=\lambda\exp\left\{-\lambda\sqrt{-\log\left(1-\rho^{2}\right)}\right\}\frac{\rho}{\left(1-\rho^{2}\right)\sqrt{-\log\left(1-\rho^{2}\right)}},$ (28)
@@
The proof is provided in Appendix A.3, and the density and log-log density plots of this PC prior are shown in Figure 9. The density plot indicates that higher values of $\lambda$ result in greater prior density at $\rho=0$, and the prior density increases sharply as $\rho$ approaches 1. In the log-log scale, the prior behavior reveals a consistent trend across different values of $\lambda$, suggesting similar characteristics despite variations in $\lambda$.
Figure 9: PC prior (for $\rho$) density (left) and log-log density (right).
@@ 4 Simulation Study and Comparisons
The logic behind our proposed prior selection framework lies in checking whether a prior favors simpler models to prevent overfitting. To further illustrate the way of using our framework, and to demonstrate the advantages of employing PC priors for circular models, this section compares the proposed priors in Section 3.2.1 with existing priors from the literature. The comparison focuses on their behavior in both the parameter scale and the distance scale (Section 4.1). Additionally, a comprehensive set of simulation studies is presented in Section 4.2.
4.1 Properties of PC and other common priors
In this section we illustrate the behavior of the proposed PC prior as well as other priors from the literature for the three cases models under consideration.
@@ 4.1.1 Von Mises Distribution
Figure 10 illustrates the behaviors of the PC prior, the popular $\operatorname{Gamma}(1,b)$ prior, $h_{2}$ prior Equation 4, and $h_{3}$ prior Equation 4 in both the parameter scale and the distance scale. The distance measure is defined between the von Mises (vM) density and the circular uniform density. For fairness, the scaling parameter $\lambda$ of the PC prior and the rate parameter $b$ of the $\operatorname{Gamma}(1,b)$ prior are calibrated so that both priors share the same median.
Figure 10: Priors for $\kappa$ in parameter (left) and distance scale (right) with the base model at $\kappa_{0}=0$. The solid line is the $\operatorname{PC}\left(\lambda=0.92\right)$ prior, the dashed line is the $\operatorname{Gamma}\left(1,0.34\right)$, the dotted line is the $h_{2}$ prior and the dot-dash line is the $h_{3}$ prior.
From the plot of the density on the distance scale, it is evident that the $h_{3}$ prior assigns zero density at $d=0$, indicating that it never suggests the data being uniform. This behavior leads to overfitting issues. The $\operatorname{Gamma}(1,b)$ prior has reasonable density at the base model, whereas it exhibits a non-monotonic density with a global maximum around $d\approx 1$ rather than at $d=0$, which indicates poor robustness in preventing overfitting. The $h_{2}$ prior does not penalize the density exponentially with respect to distance by nature; however, it has a reasonable behavior with respect to distance, that is, its density peaks at $d=0$ and decreases monotonically as $d\to\infty$.
@@
Recall that the $h_{2}$ and $h_{3}$ priors are constructed based on the MML criterion, which is intended to penalize model complexity. From the plots given above, we can see that the MML-based priors assign reasonable density to small distances from the base models but do not always allow the model to reach the boundary when the distance measure is defined by Equation 8.
@@ 4.1.2 Cardioid Distribution
Figure 13 and Figure 13 show the boundary behaviors for the priors for parameter $\ell$ with the base models being at $\ell_{0}=0$ and $\ell_{0}\rightarrow 0.5$ respectively. The plots conclude that only PC priors perform well at the boundaries, assigning non-zero density at $d=0$, whilst the density of uniform and $0.5\times\operatorname{Beta}$ priors are not even smooth. This feature illustrates the possible inherent caveat when priors are specified to be smooth and well-behaved on the parameter scale.
@@ 4.1.3 Wrapped Cauchy Distribution
The Beta prior is one of the most commonly used choices for the concentration parameter $\rho$ of the wrapped Cauchy distribution. The prior density on both the parameter scale and the distance scale is illustrated in Figure 14. The behavior of the Beta prior is particularly interesting, as it exhibits different trends near the two boundaries depending on the choice of the $a$ and $b$ hyperparameters. The right-hand-side plot shows that the Beta prior can avoid overfitting issues only when $a<1$, ensuring the density at $d=0$ is non-zero.
Figure 14: Beta prior and PC ($\lambda=1$) prior density in parameter ($p\left(\rho\right)$, left) and distance scale ($p\left(d\right)$, right).
Therefore, a Beta prior with $a<1$ can be a suitable prior choice for $\rho$. When users have sufficient information to set an informative prior, a Beta prior with $a<1$ is a reasonable choice, as it maintains the ability to prevent “overconfidence”. However, selection of the hyperparameter $b$ significantly influences the behavior of the Beta prior. Thus, we should always be careful when employing this beta prior in practice.
@@ 4.2 Simulation Study
In former subsections, we illustrate the way of employing our proposed framework, and the behavior of the discussed priors are illustrated and examined. In this section, we further show the performance of the priors and the robustness of the proposed circular PC priors through comprehensive simulation studies. The simulation study is conducted through comparing the posteriors of the concentration parameter $\xi$ ($\kappa$ for von Mises distribution, $\ell$ for cardioid distribution and $\rho$ for wrapped Cauchy distribution) under different priors. For consistency, $p\left(\mu\right)$ is chosen to be the circular uniform density, thus, the joint posterior is $p\left(\mu,\xi\mid x\right)\propto p\left(x\mid\mu,\xi\right)p\left(\mu\right)p\left(\kappa\right)$.
In the study, posterior samples are obtained using the No-U-Turn Sampler (NUTS), an adaptive variant of Hamiltonian Monte Carlo (HMC), implemented through the Stan interface (Carpenter et al., 2017; Stan Development Team, 2025). Sampling is performed using the rstan package version 2.32.7 together with Stan version 2.32.2 to ensure reproducibility. For consistency, the data are sampled from $p\left(x\mid\mu,\xi\right)$ (vM, cardioid and WC distribution respectively) with $N=100,300,1000$ sample sizes and location parameter $\mu=\pi$. In addition, the result is obtained by averaging over 100 replicates of the experiment. For reproducibility, the random seed is fixed at 520 across all experiments, with individual seeds for the 100 replicates ranging sequentially from 521 to 620. The code used in the simulation study is available at https://github.com/XiangYEstats/PCpriors-circular.
The convergence behavior of the model is assessed using the Gelman–Rubin convergence diagnostic, $\hat{R}$ (Gelman and Rubin, 1992) (see details in the Posterior Analysis section of Stan Development Team 2025). An $\hat{R}$ value close to 1 indicates good model convergence. Models with $\hat{R}>1.01$ raise convergence concerns, and $\hat{R}>1.1$ is generally considered unacceptable.
@@ 4.2.1 Von Mises Distribution
For $p\left(\kappa\right)$ we consider the $\operatorname{Gamma}\left(1,b\right)$ (exponential) priors, the $h_{2}$ and the $h_{3}$ prior (Equation 4, MML priors), and the PC priors with uniform base model (PCU, Equation 12) and point mass base model (PCP, Equation 14). The data for the simulation study are sampled with: $\kappa=0.02,1,7,59$; The hyperparameter for $\operatorname{Gamma}\left(1,b\right)$ prior has values of $b=0.01,0.05,0.1,1,5$; The scaling parameters for both PCU and PCP priors are set with $U=\frac{\pi}{2}$, and $\alpha=0.01,0.1,0.3,0.5,0.7,0.9,0.99$.
Figure 15: $\hat{R}$ values for $\kappa$ in the simulation study. The left plot shows the maximum $\hat{R}$ for each setting, and the right plot shows the $\hat{R}$ density by prior. For each setting, the $\hat{R}$ values for PCU, PCP, and exponential priors are averaged over scaling parameter values, and the $h_{2}$ and $h_{3}$ priors are averaged and presented under “MML”.
The $\hat{R}$ values for this simulation study are shown in Figure 15. The plots indicate that all priors exhibit good convergence behavior when using the Stan model, with the maximum $\hat{R}$ value being less than 1.004.
From the results in Figure 16, we can see that both PCU and PCP priors perform robustly well with different values of $\alpha$ (resulting in different values of $\lambda$), whilst the posterior for $\operatorname{Gamma}\left(1,b\right)$ priors act differently especially when the sample size is small and the concentration is high. This fact indicates that the $\operatorname{Gamma}\left(1,b\right)$ prior highly relies on the value of hyperparameter, wheares the PC priors perform well with different values of $U$ and $\alpha$. Notably, the $h_{2}$ prior also performs well in fitting the model. However, based on the discussion in Section 4.1.1, it is not recommended to employ $h_{2}$ prior when the data are highly concentrated, as it could overfit the model.
Figure 16: Simulation study for priors on the concentration parameter $\kappa$ of the von Mises distribution across different sample sizes and true $\kappa$ values. The horizontal dashed line indicates the true $\kappa$, with points showing posterior means and vertical bars indicating 95% posterior credible intervals for each prior class.
@@ 4.2.2 Cardioid Distribution
For $p\left(\ell\right)$ we consider the $\operatorname{Uniform}\left(0,0.5\right)$ prior (UNI), the $0.5\times\operatorname{Beta}\left(a,b\right)$ prior and the PC priors with the uniform base model (PCU, Equation 19) and the cardioid curve base model (PCC, Equation 22). The data are sampled with $\ell=0,0.01,0.3,0.49$; The hyperparameter for $0.5\times\operatorname{Beta}\left(a,b\right)$ prior has values of $a,b\in\left\{0.5,1,2,5\right\}$; The scaling parameters for both PCC and PCU priors are set with $U=0.5$, but with $\alpha=0.01$, 0.1, 0.3, 0.5, 0.7, 0.9, 0.99 for PCC prior and $\alpha=0.01$, 0.1, 0.2, 0.3, 0.4, 0.5 for PCU prior, since $\lambda$ takes very small values when $\alpha>0.5$ for PCU prior.
The $\hat{R}$ values for this simulation study are shown in the left-hand side of Figure 18. The plot shows that the $0.5\times\operatorname{Beta}(a,b)$ prior does not converge in the Stan model for large $\ell$ values and large sample sizes, whereas the $\operatorname{Uniform}(0,0.5)$, PCC, and PCU priors converge successfully.
Figure 17: Simulation study for priors on the concentration parameter $\ell$ of the cardioid distribution across different sample sizes and true $\ell$ values. The horizontal dashed line indicates the true $\ell$, with points showing posterior means and vertical bars indicating 95% posterior credible intervals for each prior class.
From the results given in Figure 17, we could observe that both PCU and PCC priors have much more stable and accurate performance than the $0.5\times\operatorname{Beta}\left(a,b\right)$ prior. Moreover, the PCU prior performs better than the PCC prior around the $\ell=0$ boundary, whilst the PCC prior naturally perform better around the $\ell\rightarrow 0.5$ boundary. This fact demonstrates the importance of base model selection when employing PC prior. In addition, the $\operatorname{Uniform}\left(0,0.5\right)$ prior performs well in our simulation settings, and it is able to recover the true value except when $\ell=0$ lies at the boundary. However, as discussed in Section 4.1.2, the $\operatorname{Uniform}\left(0,0.5\right)$ prior does not allocate sufficient density around both base models, and is therefore not recommended, since it may still lead to overfitting. It is worth nothing that almost all priors cannot recover the truth at the $\ell=0$ boundary, except for the $0.5\times\operatorname{Beta}\left(a,b\right)$ prior with $a<1$, and the PCU prior with small $\alpha$ values (when $U$ is fixed). This is not surprising since the estimate cannot go below 0 and it is almost impossible to recover the truth if the truth is right at the boundary.
Figure 18: $\hat{R}$ values for $\ell$ (left) and $\rho$ (right) in the simulation study. The plots show the maximum $\hat{R}$ for each setting For each setting, the $\hat{R}$ values for priors with scaling parameter are averaged over scaling parameter values.
@@ 4.2.3 Wrapped Cauchy Distribution
For $p\left(\rho\right)$ we consider the $\operatorname{Beta}\left(a,b\right)$ and the PC priors (Equation 28). The data are sampled with $\rho=0,0.01,0.5,0.99$; The hyperparameter for $\operatorname{Beta}\left(a,b\right)$ prior has values of $a,b\in\left\{0.5,1,2,5\right\}$; The scaling parameters for PC prior are set with $U=0.6$, and $\alpha=0.01,0.1,0.3,0.5,0.7,0.9,0.99$. The right-hand side plot in Figure 18 shows that both priors achieve good convergence, while the PC prior has slightly lower $\hat{R}$ values when the true $\rho$ is close to either boundary.
Figure 19 shows that the PC priors demonstrate more stable and accurate performance than the $\operatorname{Beta}\left(a,b\right)$ prior. However, if sufficient prior information supports the use of a $\operatorname{Beta}\left(a,b\right)$ prior, as discussed in Section 4.1.3, we suggest that $\operatorname{Beta}\left(a,b\right)$ prior with $a<1$ are appropriate choices, since they do not lead to overfitting issues. This behavior is also evident in Figure 19. For simulation setting with $\rho=0$, the $\operatorname{Beta}\left(a,b\right)$ prior with $a=0.5$ produces reasonable estimates, with the credible interval successfully recovering the boundary base model.
Figure 19: Simulation study for priors on the concentration parameter $\rho$ of the wrapped Cauchy distribution across different sample sizes and true $\rho$ values. The horizontal dashed line indicates the true $\rho$, with points showing posterior means and vertical bars indicating 95% posterior credible intervals for each prior class.
@@ 5 Application
In this section, we present a real-data case study to show the practical performance of the discussed priors. The data used in this study are the wind data (Figure 21) stored in “circular” package (Lund et al., 2017) in R, recorded by a meteorological station in a place named “Col de la Roa” in the Italian Alps. The dataset contains 310 measures of daily wind direction from January 29th, 2001 to March 31st, 2001 covering the data span from 3 a.m. to 4 a.m. of the day. As a circular analogue of normal distribution, von Mises distribution is employed for our model, and the Bayesian model is the same as expressed in Section 4.2. For comparison, the Stan model with $p\left(\kappa\right)$ being PC prior, $\operatorname{Gamma}\left(1,b\right)$ prior (exponential prior), $h_{2}$ prior and $h_{3}$ prior for $\kappa$ are constructed.
We choose $b=0.1,1,2,5,10$ for the $\operatorname{Gamma}\left(1,b\right)$ prior. From Figure 21, we can observe that this dataset is not highly concentrated. Therefore, for the PC prior, the circular uniform base model would be a more natural choice compared with the point mass base model. However, the purpose of comparison and illustration, models with PC prior with both base models are constructed. We choose the scaling parameter $\lambda$ following the procedure discussed in Section 3.1 with $U=\pi,\frac{\pi}{2},\frac{\pi}{4},\frac{\pi}{8},\frac{\pi}{16}$, indicating different radians around $x=0$. The $\alpha$ value is computed by finding ratio between the number of data lying in the chosen radian $U$ and 310 (total amount of data).
@@
Figure 21: Posterior means and credible intervals of $\kappa$.
Table 1: Model comparison metrics for von Mises priors. Lower WAIC/LOOIC/DIC and higher ELPD indicate better predictive performance.
Exp($\lambda$) MML PCP($\lambda$) PCU($\lambda$)
Metric 0.10 1.00 2.00 5.00 10.00 $h_{2}$ $h_{3}$ 0.23 0.84 2.49 5.83 13.14 0.38 0.40 0.59 0.81 1.15
WAIC 838.29 838.29 838.41 838.59 839.61 838.47 838.35 838.35 838.27 838.34 838.34 838.38 838.27 838.44 838.22 838.31 838.34
LOOIC 838.29 838.30 838.41 838.59 839.61 838.47 838.35 838.36 838.28 838.35 838.34 838.38 838.27 838.44 838.22 838.31 838.34
ELPD -419.15 -419.15 -419.20 -419.29 -419.80 -419.23 -419.17 -419.18 -419.14 -419.17 -419.17 -419.19 -419.13 -419.22 -419.11 -419.15 -419.17
DIC 836.10 836.13 836.22 836.51 837.59 836.19 836.14 836.14 836.11 836.13 836.12 836.15 836.11 836.18 836.09 836.13 836.14
After fitting the model, the posterior mean and the 95% credible interval for $\kappa$ are given in Figure 21. In the plot, the horizontal dashed line indicates the frequentist maximum likelihood estimate of $\kappa$, with points showing posterior means and vertical bars indicating 95% posterior credible intervals for each prior class. From the plot, the PC prior with circular uniform base model (PCU) and PC prior with point mass base model (PCP) fit the data well in terms of the stability of credible interval and posterior mean. For this dataset, as a more natural choice, the PCU prior is indeed robust than PCP prior, since the value of scaling parameter $\lambda$ does not vary much with the same given $U$ and $\alpha$ values.
The model comparison metrics are given in Table 1. The Watanabe–Akaike Information Criterion (WAIC) (Watanabe and Opper, 2010), Leave-One-Out Information Criterion (LOOIC) (Vehtari et al., 2017), Expected Log Predictive Density (ELPD) and Deviance Information Criterion (DIC) (Spiegelhalter et al., 2002) are employed for assessing the prior performance in same model. The $\operatorname{PCU}\left(0.59\right)$ prior performs the best among all priors according to the score of all metrics. The other PCU and PCP priors show robustness, compared with the $\operatorname{Gamma}\left(1,b\right)$ (Exp) prior, as the most extreme values for $\lambda$ in our example still lead to good model comparison metrics values, whereas Figure 21 shows that the posterior mean and credible interval of Exp prior is quite sensitive to the $\lambda$ values, but there is insufficient guidance for choosing the values for hyperparameters of $\operatorname{Gamma}\left(1,b\right)$ prior. In addition, the $h_{2}$ and $h_{3}$ priors both perform well in this example. In Section 4.1.1, we discussed that $h_{2}$ prior also performs well in preventing overfitting when the base model is the circular uniform distribution. Thus, to prevent overfitting, the $h_{2}$ prior also appears to be a proper choice of prior for $\kappa$ when the data appear to have low concentration.